mathslover (mathslover):

integrate: (1+5t+t^2)/1

OpenStudy (unklerhaukus):

with respect to t?

OpenStudy (lgbasallote):

over 1?

mathslover (mathslover):

yes ....

OpenStudy (lgbasallote):

\[\int (1 + 5t + t^2)dt = \int dt + \int 5tdt + \int t^2dt\]

mathslover (mathslover):